OpenStudy (anonymous):

Evaluate 9x^2y^-2 for x = –3 and y = 2

OpenStudy (anonymous):

9x^2y^-2 = 9(-3)^2(2)^-2 = -27^4^-2 = 531,441^-2 = 1/(531,442^2) = 3.54*10^-12

OpenStudy (anonymous):

A. 324 B. 20 1/4 C. 9 (-6) D. 1/ 144 which one would it be?